Can anyone help out? Julie and Eric row their boat (at a constant speed) 63 miles downstream for 7 hours, helped by the current. Rowing at the same rate, the trip back against the current takes 9 hours. Find the rate of the current.
Answer by Paul(988)

63(t-7)=63(9-t)
63t-441=567-63t
126t-441+441=567+441
126t=1008
t=8
hence the rate of the current is 8miles per hour.